# GAST: Sequential Gaussian Avatars with Hierarchical Spatio-temporal Context

### v1

@@ 11 Details of Optimization
In our experiments, the number $\lvert\mathcal{S}\rvert$ (in Eq. (16)) of sequence sampled in different scales is set to 3 and each sequence length $L$ (in Eq. (6)) is set to 8. We use the $\tau=8$ nearest SMPL template vertexes’ velocities to get each Gaussian primitive’s local velocity embedding $e_{i}$ (in Eq. (11)). Additionally, we utilize the adaptive densification in  [23] to control the number of Gaussian primitives.
For DNA-Rendering dataset, we use the SMPL-X [42] model with pose dimension $P\in\mathbb{R}^{25\times 3}$ in our experiments. We adopt the SMPL-X template with $N=10475$ vertexes as Gaussian initialization. We optimize 30k iterations and set the loss weights $\lambda_{0}=1.0,\lambda_{1}=0.01,\lambda_{2}=0.01$. The initial sampling step and the increasing interval are $s_{0}=1$ and $\Delta s=2$, respectively. The final sample scales are $\mathcal{S}=\{1,3,5\}$ (in Eq. (16)). For I3D-Human dataset, we follow Dyco [5] to use SMPL [37] model in our experiments for fairness. The number of SMPL template vertexes used for Gaussian initialization is $N=6890$ and the corresponding pose dimension is $P\in\mathbb{R}^{23\times 3}$. We optimize 15k iterations and set the loss weights $\lambda_{0}=1.0,\lambda_{1}=0.1,\lambda_{2}=0.1$. We set the increasing interval $\Delta s=9$. The initial sampling step is $s_{0}=24$ and the final sample scales are $\mathcal{S}=\{24,33,42\}$ (in Eq. (16)). Note that the experiment in Tab. 3 (d) adopts single sampling step that $\mathcal{S}=\{s_{0}\}$.
@@ 12 Additional Results
12.1 Comparisons of Novel View Rendering
Figure 7: Additional Novel Pose Qualitative Results on I3D-Human Dataset [5].
@@ 12.1 Comparisons of Novel View Rendering
12.2 Comparisons of Novel Pose Rendering
Table 4: Quantitative Results of Novel Pose Renderings on I3D-Human Dataset.
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
3DGS-Avatar [44] 29.53 0.9554 38.58
Dyco [5] 30.30 0.9582 33.52
GauHuman [17] 29.31 0.9525 48.29
GAST(Ours) 30.39 0.9582 38.15
Tab. 4 reports the quantitative results of novel pose rendering on the I3D-Human dataset [5]. Our method achieves competitive performance compared with SOTA human modeling methods. Fig. 7 shows the qualitative comparisons of novel pose rendering and we utilize error maps to visualize the differences between various methods more clearly. Compared to Dyco [5], which relies solely on joint motions as conditions, our GAST leverages the motion state of each Gaussian primitive to predict deformations, enabling more flexible modeling of motions in regions distant from the human body. The shape of the clothing in Fig. 7 demonstrates that our method can render the appearance consistent with the novel pose motion, leading to more accurate results.
@@ 13 Additional Ablations
$\mathcal{S}=\{24\}$ 32.05 0.96468 32.82
$\mathcal{S}=\{33\}$ 32.04 0.96467 33.06
$\mathcal{S}=\{42\}$ 32.06 0.96472 32.87
$\mathcal{S}=\{24,33,42\}$ 32.13 0.96476 32.65
@@
Table 6: Impact of Different KNN Vertex Number $\tau$.
KNN num. $\tau$ $\tau=1$ $\tau=3$ $\tau=6$ $\tau=8$
PSNR $\uparrow$ 32.06 32.07 32.08 32.13

### v2

@@ 11 Details of Optimization
In our experiments, the number $\lvert\mathcal{S}\rvert$ (in Eq. (14)) of sequence sampled in different scales is set to 3, and each sequence length $L$ (in Eq. (6)) is set to 8. We use the $\tau=8$ nearest SMPL template vertexes’ velocities to get each Gaussian primitive’s local velocity embedding $e_{i}$ (in Eq. (12)). Additionally, we utilize the adaptive densification in  [27] to control the number of Gaussian primitives.
For the DNA-Rendering dataset, we use the SMPL-X [51] model with pose dimension $P\in\mathbb{R}^{25\times 3}$ in our experiments. We adopt the SMPL-X template with $N=10475$ vertexes as Gaussian initialization. We optimize 30k iterations and set the loss weights $\lambda_{0}=1.0,\lambda_{1}=0.01,\lambda_{2}=0.01$. The initial sampling step and the increasing interval are $s_{0}=1$ and $\Delta s=2$, respectively. The final sample scales are $\mathcal{S}=\{1,3,5\}$ (in Eq. (14)). For the I3D-Human dataset, we follow Dyco [8] to use the SMPL [43] model in our experiments for fairness. The number of SMPL template vertexes used for Gaussian initialization is $N=6890$ and the corresponding pose dimension is $P\in\mathbb{R}^{23\times 3}$. We optimize 15k iterations and set the loss weights $\lambda_{0}=1.0,\lambda_{1}=0.1,\lambda_{2}=0.1$. We set the increasing interval $\Delta s=9$. The initial sampling step is $s_{0}=24$ and the final sample scales are $\mathcal{S}=\{24,33,42\}$ (in Eq. (14)). Note that the experiment in Tab. 4 (d) adopts single sampling step that $\mathcal{S}=\{s_{0}\}$.
Figure 8: Novel View Qualitative Results on I3D-Human.
@@ 12 Additional Results
12.1 Experiments on ZJU-MoCap
ZJU-MoCap [53] is a common benchmark dataset in human avatar, which is mainly collected under controlled speeds and tight-fitting garments. We use 6 sequences (377, 386, 387, 392, 393, 394) for experiments following the dataset split in 3DGS-Avaar[56].
Comparison on ZJU-MoCap. Results on DNA-Rendering and I3D-Human demonstrate that our methods achieve better performance in complex cases with loose clothing and uncontrolled speeds. Furthermore, Tab. 5 and Fig. 9 show that our method also achieves competitive metrics in simpler controlled scenarios, which verifies the generalizability of the proposed methods.
Table 5: Quantitative Results on ZJU-MoCap.
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
Dyco [8] 30.37 0.9599 29.47
3DGS-Avatar [56] 30.59 0.9609 27.09
GauHuman [22] 31.04 0.9620 31.81
Ours 31.02 0.9619 28.89
Figure 9: Qualitative Results on ZJU-MoCap.
12.2 Visual Comparisons On DNA-Rendering
@@
12.3 Video Quality
We provide rendered videos on the DNA-Rendering dataset [9], along with the corresponding error maps similar to Fig. 3. Please refer to the attached videos for comparison. In the videos, the novel view results are rendered on temporal frames with pose IDs 0–99, using camera viewpoints that are held out during training.
@@ 13 Additional Ablations
$\mathcal{S}=\{24\}$ 32.01 0.9655 30.71
$\mathcal{S}=\{33\}$ 32.09 0.9656 30.88
$\mathcal{S}=\{42\}$ 32.11 0.9658 30.81
$\mathcal{S}=\{24,33,42\}$ 32.24 0.9664 29.78
@@
Table 7: Impact of Different KNN Vertexe Number $\tau$.
KNN num. $\tau$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
$\tau=1$ 32.14 0.9659 30.41
$\tau=5$ 32.15 0.9660 30.32
$\tau=8$ 32.24 0.9664 29.79
$\tau=10$ 32.20 0.9663 30.11
14 Computational Cost Analysis
Table 8: Computational Efficiency.
Mehods Train Time FPS Train Mem.
Dyco 6 h 0.7 20 GB
3DGS-Avatar 25 m 25 6 GB
Ours ($3k$ $iter$) 5 m 62 8 GB
We provide computational cost analysis for experiments on ZJU-Mocap (conducted on a single 4090 GPU). Tab. 8 shows that our approach achieves competitive computational efficiency. The additional computation time for each frame’s Vertex Motion Template (multi-scale sampling number $|\mathcal{S}|=3$, sequence length $L=8$) is about 0.09 s, and it only needs to be computed once during preprocessing before training.